# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ Abstract.
We study the sine–Gordon measure defined on each homotopy class. The energy space decomposes into infinitely many such classes indexed by the topological degree $Q\in\mathbb{Z}$. Even though the sine–Gordon action admits no minimizer in homotopy classes with $|Q|\geq 2$, we prove that the Gibbs measure on each class nevertheless concentrates and exhibits Ornstein–Uhlenbeck fluctuations near the multi-soliton manifold in the joint low-temperature and infinite-volume limit. Furthermore, we show that the joint distribution of the multi-soliton centers coincides with the ordered statistics of independent uniform random variables, so that each soliton’s location follows a Beta distribution.
@@
topological solitons; multi-soliton manifold; Ornstein–Uhlenbeck fluctuations, Large deviations, expected location, expected gap
@@ 1.1. Motivations and implications of the main results
$E(\phi)=\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\phi|^{2}dx+\int_{\mathbb{R}}(1-\cos\phi)dx.$ (1.1)
@@
$Q=\frac{1}{2\pi}(\phi(\infty)-\phi(-\infty))\in\mathbb{Z}.$ (1.2)
@@
$\mathcal{C}_{Q}=\{\phi\in\mathcal{C}:(\phi(\infty)-\phi(-\infty))/2\pi=Q\}.$ (1.3)
@@
    $|Q|\geq 2$ (higher-charge sectors): no minimizer exists. For higher charge, the lack of compactness prevents existence of a minimizer.
Although the higher-charge sectors $|Q|\geq 2$ admit no minimizer, we nevertheless study, for each $Q\in\mathbb{Z}$, the concentration and fluctuation behavior of the Gibbs measure [^1] Here $Z_{\varepsilon}$ denotes partition functions that may differ from line to line. $\rho_{\varepsilon}^{Q}$ on the homotopy class $\mathcal{C}_{Q}$
$\rho_{\varepsilon}^{Q}(d\phi)=Z_{\varepsilon}^{-1}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi(x))dx-\frac{1}{2\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}d\phi(x)$
@@
    $\min_{i\neq j}|\xi_{j}-\xi_{i}|\geq\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|\to\infty$
@@
    At leading order, typical configurations under the Gibbs ensemble consist of exactly $|Q|$ solitons whose mutual separations are of order $\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$. Consequently, configurations in which the solitons collide are unlikely, and the solitons behave as effectively non-interacting objects.
@@ 1.2. Main results
$\mu_{\varepsilon}^{Q}(d\phi)=\frac{1}{Z_{\varepsilon}^{Q}}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}d\phi(x),$ (1.4)
@@
$\rho_{\varepsilon}^{Q}(d\phi)=Z_{\varepsilon}^{-1}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi(x))dx\Big\}\mu_{\varepsilon}^{Q}(d\phi).$ (1.5)
@@
    $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-c\delta^{2},$ (1.6)
@@
    $\mathcal{M}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\Big\},\quad Q>0.$ (1.7)
@@
    $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-ce^{-d},$ (1.8)
    for any $\delta>0$, $d>0$ satisfying [^3] Later, we prove that the collision manifold $\mathcal{M}_{Q}^{<d}$ is an unlikely event. Therefore, when the field $\phi$ is sufficiently close to $\mathcal{M}_{Q}^{<d}$, in the sense that $\frac{1}{2c}e^{-d}\geq\delta$, we obtain the same result. See Lemmas 3.10 and 3.12. $\frac{1}{2c}e^{-d}\geq\delta$, where the collison manifold $\mathcal{M}_{Q}^{<d}$, $Q>0$, is defined as
    $\mathcal{M}^{<d}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|<d\Big\}.$ (1.9)
@@
The first part of Theorem 1.2 shows that when a field $\phi$ is far from the multi-soliton manifold $\mathcal{M}_{Q}$, the event is exponentially unlikely with rate $\delta^{2}$. In particular, the manifold consists of exactly $|Q|$ solitons, without any mixture of kinks and anti-kinks. Furthermore, the second part shows that although a field $\phi$ is close to the multi-soliton manifold, when the solitons collide so that $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$, the collision region becomes a large-deviation event. From (1.6), (1.8), and the condition $\frac{1}{2c}e^{-d}\geq\delta$, we may choose the distance and collision scales as $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$. With these choices, we obtain
@@
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\leq e^{-c\varepsilon^{-\frac{1}{2}+}}$
for some $c>0$ as $\varepsilon\to 0$. An interesting fact is that we identify the collision scale $\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$ that determines the typical behavior of solitons. Hence, most of the probability mass is concentrated in the well-separated (non-collision) region
@@
In the proof of Theorem 1.2, the key aspect is to understand how the solitons interact with each other on the collision scale $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$, and how the energy behaves even though these configurations are not minimizers.
@@
In Theorem 1.2, there is a competition between the vanishing energy scale $\varepsilon\to 0$ and the entropic effects arising from the growing interval $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}\to\infty$. Under our method, the interval size $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ is optimal. See Remark 6.2 for an explanation of this scaling.
As a consequence of Theorem 1.2, under the measure $\rho^{Q}_{\varepsilon}$, the leading-order behavior is described by $\sum_{j=1}^{Q}m(\cdot-\xi_{j})$ with almost no collision $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}|\to\infty$. In the following theorem, we investigate the next-order fluctuation behavior around the multi-solitons. To state our next theorem, we first introduce $\pi^{\varepsilon}$, the projection onto the (approximating) [^4] Since we work on the finite volume $L_{\varepsilon},L_{\varepsilon}]$, we need to define an approximating multi–soliton profile $\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})$ so that the entire transition from 0 to $2\pi Q$ occurs inside this interval. As $\varepsilon\to 0$, $m^{\varepsilon}(\cdot-\xi_{j})$ becomes a more and more precise approximation of the topological soliton $m(\cdot-\xi_{j})$ on $\mathbb{R}$ (see (4.3)) multi-soliton manifold $\mathcal{M}_{Q}^{\varepsilon,\geq(1+\eta)d_{\varepsilon}}$ defined in (4.15), where the measure $\rho^{Q}_{\varepsilon}$ concentrates (see Remarks 6.4, 6.5, and (6.32)).
@@
under the separation scale $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$ (see also (5.1)). It allows to study the Gaussian measure $\exp\big\{-\frac{1}{2}\langle\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})v,v\rangle\big\}$ whose covariance structure (Lemma 9.3) and correlation decay (Proposition 8.3) are crucial ingredients in the proof of the central limit theorem.
@@
Before stating the final theorem, we first present some preliminaries. Under the coordinate representation [^5] Thanks to the large deviation results in Theorem 1.2, we can write the field $\phi$ as a multi–soliton configuration plus a small perturbation. $\phi=\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})+v$, where $\|v\|_{L^{2}}<\delta$, defined in (4.15), we consider the projection $\pi_{\varepsilon}^{T}(\phi)=(\xi_{1},\dots,\xi_{Q})$ onto the coordinate variables, where666when working on the finite volume $L_{\varepsilon},L_{\varepsilon}]$, we need to define an approximating multi–soliton profile $\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})$ so that the entire transition from 0 to $2\pi Q$ occurs inside this interval. For this reason, we introduce a slightly smaller interval, denoted by $\overline{L}_{\varepsilon}\sim L_{\varepsilon}$, on which the translation $\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ modes are defined (see (4.2)). $\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}$. Furthermore, we define the (marginal) tangential projection $\pi^{T}_{j}(\phi)=\xi_{j}$, where $\xi_{j}$ denotes the $j$-th ordered center in the increasing rearrangement $\xi_{1}\leq\cdots\leq\xi_{Q}$.
@@ 1.3.3. Discrete focusing NLS Gibbs measure and single-soliton manifold
The discretized focusing Gibbs measure and the nonlinear Schrödinger (NLS) equation on a lattice have been studied. When $d\geq 3$, Chatterjee–Kirkpatrick [12] initiated the study of the discretized focusing NLS Gibbs measure, identified a critical temperature, and showed that below this threshold the system exhibits striking single-soliton–like behavior. In [13] Chatterjee used microcanonical invariant measures and showed that a typical function in the ensemble decomposes into a “visible” part, which is close to a single soliton, and an “invisible” part that is small in the $L^{\infty}$ norm. Notice that, in particular, regarding the reason why a single-soliton profile appears, Chatterjee mentioned in [13] that “ whereas multisoliton solutions eventually merge into a single soliton on the finite discrete torus considered in Theorem 1”. In contrast to that situation, in our sine–Gordon model, the boundary conditions imposed by a fixed homotopy class (1.3) enforce the presence of multiple solitons, and therefore prevent the multi-soliton configuration from collapsing into a single soliton. Again, as emphasized above, to the best of our knowledge, our results are the first study of the concentration and fluctuations of Gibbs measures around multi-soliton manifold. In particular, a crucial aspect is to understand how solitons interact with each other on the collision scale $\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$, and how the energy behaves at this scale, even though such configurations are not minimizers.

### v2

@@ Abstract.
We study the sine–Gordon measure defined on each homotopy class. The energy space decomposes into infinitely many such classes indexed by the topological degree $Q\in\mathbb{Z}$. Even though the sine–Gordon action admits no minimizer in homotopy classes with $|Q|\geq 2$, we prove that the Gibbs measure on each class nevertheless concentrates and exhibits Ornstein–Uhlenbeck fluctuations near the multi-soliton manifold in the joint low-temperature and infinite-volume limit. Moreover, we show that soliton collisions are unlikely events, so that typical states consist of solitons separated at an appropriate scale. Finally, we identify the joint distribution of the multi-soliton centers as the ordered statistics of independent uniform random variables, so that each soliton’s location follows a Beta distribution.
@@
topological solitons; multi-soliton manifold; collision manifold; large deviations, Ornstein–Uhlenbeck fluctuations, expected location, expected gap
@@ 1.1. Motivations and implications of the main results
(1.1) $E(\phi)=\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\phi|^{2}dx+\int_{\mathbb{R}}(1-\cos\phi)dx.$
@@
(1.2) $Q=\frac{1}{2\pi}(\phi(\infty)-\phi(-\infty))\in\mathbb{Z}.$
@@
(1.3) $\mathcal{C}_{Q}=\{\phi\in\mathcal{C}:(\phi(\infty)-\phi(-\infty))/2\pi=Q\}.$
@@
    $|Q|\geq 2$ (higher-charge sectors): no minimizer exists. For higher charge, the lack of compactness prevents the existence of a minimizer: the energy infimum is approached only by a “runaway” configuration of $|Q|$ widely separated kinks/antikinks.
Although the higher-charge sectors $|Q|\geq 2$ admit no minimizer, we nevertheless study, for each $Q\in\mathbb{Z}$, the concentration and fluctuation behavior of the Gibbs measure $\rho_{\varepsilon}^{Q}$ on the homotopy class $\mathcal{C}_{Q}$
$\rho_{\varepsilon}^{Q}(d\phi)=(Z_{\varepsilon}^{Q})^{-1}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi(x))dx-\frac{1}{2\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}d\phi(x)$
@@
    $\min_{i\neq j}|\xi_{j}-\xi_{i}|\geq\big|\log(\varepsilon\log\tfrac{1}{\varepsilon})\big|\to\infty$
@@
    At leading order, typical configurations under the Gibbs ensemble consist of exactly $|Q|$ solitons whose mutual separations are of order $\big|\log(\varepsilon\log\tfrac{1}{\varepsilon})\big|$. Consequently, configurations in which the solitons collide are unlikely, and the solitons behave as effectively non-interacting objects.
@@ 1.2. Main results
(1.4) $\mu_{\varepsilon}^{Q}(d\phi)=\frac{1}{Z_{\varepsilon}^{Q,\textup{BB}}}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}d\phi(x),$
@@
(1.5) $\rho_{\varepsilon}^{Q}(d\phi)=(Z_{\varepsilon}^{Q})^{-1}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi(x))dx\Big\}\mu_{\varepsilon}^{Q}(d\phi).$
@@
    (1.6) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-c\delta^{2},$
@@
    (1.7) $\mathcal{M}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\Big\},\quad Q>0.$
@@
    (1.8) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-ce^{-d},$
    for any $\delta>0$, $d>0$ satisfying [^2] Later, we prove that the collision manifold $\mathcal{M}_{Q}^{<d}$ is an unlikely event. Therefore, when the field $\phi$ is sufficiently close to $\mathcal{M}_{Q}^{<d}$, in the sense that $ce^{-d}\geq\delta^{2}$, we obtain the same result. See Lemmas 3.10 and 3.12. $ce^{-d}\geq\delta^{2}$, where the collision manifold $\mathcal{M}_{Q}^{<d}$, $Q>0$, is defined as
    (1.9) $\mathcal{M}^{<d}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|<d\Big\}.$
@@
The first part of Theorem 1.2 shows that when a field $\phi$ is far from the multi-soliton manifold $\mathcal{M}_{Q}$, the event is exponentially unlikely with rate $\delta^{2}$. In particular, the manifold consists of exactly $|Q|$ solitons, without any mixture of kinks and anti-kinks. Furthermore, the second part shows that although a field $\phi$ is close to the multi-soliton manifold, when the solitons collide so that $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$, the collision region becomes a large-deviation event. From (1.6), (1.8), and the condition $ce^{-d}\geq\delta^{2}$, we may choose the distance and collision scales as $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$. With these choices, we obtain
@@
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\leq e^{-c\log\frac{1}{\varepsilon}}$
for some $c>0$ as $\varepsilon\to 0$. An interesting fact is that we identify the collision scale $\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$ that determines the typical behavior of solitons. Hence, most of the probability mass is concentrated in the well-separated (non-collision) region
@@
In the proof of Theorem 1.2, the key aspect is to understand how the solitons interact with each other on the collision scale $d_{\varepsilon}=\big|\log(\varepsilon\log\tfrac{1}{\varepsilon})\big|$, and how the energy behaves even though these configurations are not minimizers.
@@
In Theorem 1.2, there is a competition between the vanishing energy scale $\varepsilon\to 0$ and the entropic effects arising from the growing interval $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}\to\infty$. Under our method, the interval size $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ is optimal. See Section 6 for an explanation of this scaling.
As a consequence of Theorem 1.2, under the measure $\rho^{Q}_{\varepsilon}$, the leading-order behavior is described by $\sum_{j=1}^{Q}m(\cdot-\xi_{j})$ with almost no collision $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq|\log(\varepsilon\log\tfrac{1}{\varepsilon})|\to\infty$. In the following theorem, we investigate the next-order fluctuation behavior around the multi-solitons. To state our next theorem, we first introduce $\pi^{\varepsilon}$, the projection onto the (approximating) [^3] Since we work on the finite volume $L_{\varepsilon},L_{\varepsilon}]$, we need to define an approximating multi–soliton profile $\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})$ so that the entire transition from 0 to $2\pi Q$ occurs inside this interval. As $\varepsilon\to 0$, $m^{\varepsilon}(\cdot-\xi_{j})$ becomes a more and more precise approximation of the topological soliton $m(\cdot-\xi_{j})$ on $\mathbb{R}$ (see (4.3)) multi-soliton manifold $\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}}$ defined in (4.15), where the measure $\rho^{Q}_{\varepsilon}$ concentrates (see Lemma 7.1).
@@
under the separation scale $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$ (see also (5.1)). It allows to study the Gaussian measure $\exp\big\{-\frac{1}{2}\langle\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})v,v\rangle\big\}$ whose covariance structure (Lemma 9.3) and correlation decay (Proposition 8.3) are crucial ingredients in the proof of the central limit theorem.
@@
Before stating the final theorem, we first present some preliminaries. Under the coordinate representation [^4] Thanks to the large deviation results in Theorem 1.2, we can write the field $\phi$ as a multi–soliton configuration plus a small perturbation. $\phi=\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})+v$, where $\|v\|_{L^{2}}<\delta$, defined in (4.15), we consider the projection $\pi_{\varepsilon}^{T}(\phi)=(\xi_{1},\dots,\xi_{Q})$ onto the coordinate variables, where555when working on the finite volume $L_{\varepsilon},L_{\varepsilon}]$, we need to define an approximating multi–soliton profile $\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})$ so that the entire transition from 0 to $2\pi Q$ occurs inside this interval. For this reason, we introduce a slightly smaller interval, denoted by $\overline{L}_{\varepsilon}\sim L_{\varepsilon}$, on which the translation $\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ modes are defined (see (4.2)). $\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}$. Furthermore, we define the (marginal) tangential projection $\pi^{T}_{j}(\phi)=\xi_{j}$, where $\xi_{j}$ denotes the $j$-th ordered center in the increasing rearrangement $\xi_{1}\leq\cdots\leq\xi_{Q}$.
@@ 1.3.3. Discrete focusing NLS Gibbs measure and single-soliton manifold
The discretized focusing Gibbs measure and the nonlinear Schrödinger (NLS) equation on a lattice have been studied. When $d\geq 3$, Chatterjee–Kirkpatrick [12] initiated the study of the discretized focusing NLS Gibbs measure, identified a critical temperature, and showed that below this threshold the system exhibits striking single-soliton–like behavior. In [13] Chatterjee used microcanonical invariant measures and showed that a typical function in the ensemble decomposes into a “visible” part, which is close to a single soliton, and an “invisible” part that is small in the $L^{\infty}$ norm. Notice that, in particular, regarding the reason why a single-soliton profile appears, Chatterjee mentioned in [13] that “ whereas multisoliton solutions eventually merge into a single soliton on the finite discrete torus considered in Theorem 1”. In contrast to that situation, in our sine–Gordon model, the boundary conditions imposed by a fixed homotopy class (1.3) enforce the presence of multiple solitons, and therefore prevent the multi-soliton configuration from collapsing into a single soliton. Again, as emphasized above, to the best of our knowledge, our results are the first study of the concentration and fluctuations of Gibbs measures around multi-soliton manifold. In particular, a crucial aspect is to understand how solitons interact with each other on the collision scale $\big|\log(\varepsilon\log\tfrac{1}{\varepsilon})\big|$, and how the energy behaves at this scale, even though such configurations are not minimizers.